OpenStudy (anonymous):

Ball 1, with a mass of 10kg, is moving to the right at 20 m/s. Ball 2 is moving to the left at 20 m/s. Upon collision, ball 2 comes to a standstill, and ball 1 moves to the left at twice its origional speed. What is the mass of ball 2?

OpenStudy (anonymous):

 Momentum is conserved: momentum before collision = momentum after collision \[m _{1}v _{1} + m _{2}v _{2} = m _{1}v _{1} + m _{2}v _{2}\] \[(10 \times 20) + (m \times -20) = (10 \times -40) + (m \times 0)\] \[200 - 20m = -400\] \[m = (-400 - 200)/(-20) = 30 kg\]